# Deterministic Dynamic Maximal Matching in Sublinear Update Time

### v1

@@ 6.3.3 Initializing the matchings $M_{\texttt{hilo}}$ and $M_{\texttt{base}}$ at the start of a phase
Throughout this section, we focus on the snapshort of time at the start of a concerned phase, just before the first update (in that phase) arrives. Thus, for example, throughout this section we have $V_{\texttt{sf}}=V_{\texttt{hi}}$ and $E(H_{\texttt{init}})=E(H_{\texttt{core}})$ (see Definition 6.2 and Corollary 6.3). Furthermore, recall that $M(S)$ is the set of edges in a matching $M$ that are incident on $S$ (see Section 2).
@@
    The total time spent to compute $M_{\texttt{hilo}}$ and $M_{\texttt{base}}$ is proportional to: $O(nB)$, plus the time it takes to compute $M_{\texttt{most}}$ (see Claim 6.7), plus the time it takes to make one call to $\mathrm{Init}(.)$ and $6\delta n$ calls to $\mathrm{Augment}(.)$ in $H_{\texttt{hilo}}$ (see Lemma 5.1 and part-(1) of Claim 6.8).
Our algorithm will satisfy the following invariant, which implies part-(1) of Claim 6.9.
@@
We devote the rest of this section to the proof of Claim 6.9. Since $M_{\texttt{most}}\left(V_{\texttt{sf}}\right)\subseteq M_{\texttt{most}}\subseteq E(H_{\texttt{init}})=E(H_{\texttt{core}})$ (see Claim 6.7), from (8) it follows that:
@@
Fix the parameters $\gamma,X$ as in (9), and recall Claim 6.8. We will run the LPM data structure on $H_{\texttt{hilo}}$ (see Lemma 5.1), and we will let $M_{\texttt{hilo}}$ be the matching $M$ maintained by LPM.
{mdframed}
[hidealllines=true, backgroundcolor=gray!20, leftmargin=0cm,innerleftmargin=0.4cm,innerrightmargin=0.4cm,innertopmargin=0.4cm,innerbottommargin=0.4cm,roundcorner=10pt] We first set $M_{\texttt{base}}\leftarrow M_{\texttt{most}}$, and then call $\mathrm{Init}(M_{\texttt{base}}(V_{\texttt{sf}}))$ in LPM. [^7] This is a valid operation, since $M_{\texttt{most}}(V_{\texttt{sf}})\subseteq E(H_{\texttt{hilo}})$ according to (11).
@@
At the end of the above call to $\mathrm{Init}(.)$, at most $6\delta n$ nodes from $V_{\texttt{sf}}\cup V_{\texttt{med}}$ are free under $M_{\texttt{base}}$, and 6.10 is satisfied.
@@
Follows from Claims 6.7 and 6.8 and Lemma 5.1. ∎
{mdframed}
[hidealllines=true, backgroundcolor=gray!20, leftmargin=0cm,innerleftmargin=0.4cm,innerrightmargin=0.4cm,innertopmargin=0.4cm,innerbottommargin=0.4cm,roundcorner=10pt] Let $F^{\star}=V_{\texttt{sf}}\setminus V(M_{\texttt{hilo}})$ denote the set of safe high nodes that are free under $M_{\texttt{hilo}}$ at this point. By Corollary 6.11 and Claim 6.8, we have $|F^{\star}|\leq 6\delta n$, and $\mathrm{deg}_{H_{\texttt{hilo}}}(v)\geq X$ for all $v\in F^{\star}$. Next, for each node $v\in F^{\star}$, we call the subroutine $\mathrm{MatchViaAugment}(v)$, as described below. This concludes our initialization of the matching $M_{\texttt{base}}$ at the start of a phase.
{mdframed}
[hidealllines=true, backgroundcolor=gray!20, leftmargin=0cm,innerleftmargin=0.4cm,innerrightmargin=0.4cm,innertopmargin=0.4cm,innerbottommargin=0.4cm,roundcorner=10pt]
@@
We call this subroutine only if $v$ is currently unmatched in $M_{\texttt{hilo}}$ and $v\in V_{\texttt{sf}}$, so that $v$ lies on the “left” with $\mathrm{deg}_{\texttt{hilo}}(v)\geq X$ (see Claim 6.8). It performs the following operations.
  • Make a call to $\mathrm{Augment}(v)$ in LPM (see Lemma 5.1). This step matches $v$ under $M_{\texttt{hilo}}$ by applying an augmenting path, adds exactly one node (say, $u_{v}$) besides $v$ to the set $V(M_{\texttt{hilo}})$, and does not remove any node from the set $V(M_{\texttt{hilo}})$. Since $v$ lies to the “left” of the bipartite graph $H_{\texttt{hilo}}$, it follows that $u_{v}$ lies to the “right” of $H_{\texttt{hilo}}$ and hence $u_{v}\in V_{\texttt{lo}}\cup V_{\texttt{a-lo}}$ (see Claim 6.8). Let $M^{\texttt{old}}_{v}$ (resp. $M^{\texttt{new}}_{v}$) denote the set of edges that get removed from (resp. added to) the matching $M_{\texttt{hilo}}$ as we apply this augmenting path.
  • Set $M_{\texttt{base}}\leftarrow M_{\texttt{base}}\cup M^{\texttt{new}}_{v}\setminus M^{\texttt{old}}_{v}$.
  • If $u_{v}$ was previously matched under $M_{\texttt{base}}$ (say, via an edge $(u_{v},u^{\prime}_{v})\in M_{\texttt{base}}$), then set $M_{\texttt{base}}\leftarrow M_{\texttt{base}}\setminus\{(u_{v},u^{\prime}_{v})\}$.
@@
For part-(2), the node $v$ gets added to the set $V(M_{\texttt{base}})$ because of the call to $\mathrm{Augment}(v)$ in LPM (see Lemma 5.1), and because we ensure that $M_{\texttt{base}}\supseteq M_{\texttt{hilo}}$ (see part-(1) above). Furthermore, $u^{\prime}_{v}$ is the only node that can potentially get removed from $V(M_{\texttt{base}})$. Just before the call to $\mathrm{MatchViaAugment}(v)$, we have $(u_{v},u^{\prime}_{v})\notin M_{\texttt{hilo}}$ (otherwise the augmenting path we apply will not end at $u_{v}$) and $M_{\texttt{hilo}}=M_{\texttt{base}}\left(V_{\texttt{sf}}\right)$ (due to 6.10). Thus, it follows that $u^{\prime}_{v}\notin V_{\texttt{sf}}$.
@@
Proof of Claim 6.9
Part-(1) follows from 6.10, Corollary 6.11 and part-(1) of Claim 6.12.
For part-(2), note that $\left|\left(V_{\texttt{sf}}\cup V_{\texttt{med}}\right)\setminus V(M_{\texttt{base}})\right|\leq 6\delta n$ at the end of the call to $\mathrm{Init}(.)$ in LPM (see Corollary 6.11). Subsequently, for each node $v\in F^{\star}=V_{\texttt{sf}}\setminus V(M_{\texttt{hilo}})$, we call $\mathrm{MatchViaAugment}(v)$, which leads to:
  • the node $v$ getting added to $V(M_{\texttt{base}})$ and no node already in $V_{\texttt{sf}}\cap V(M_{\texttt{base}})$ getting deleted from $V(M_{\texttt{base}})$ (see part-(2) of Claim 6.12), and
  • at most two nodes in $V_{\texttt{med}}$ getting deleted from $V(M_{\texttt{base}})$ (see part-(3) of Claim 6.12).
@@
For part-(3), note that $H_{\texttt{hilo}}$ is a subgraph of $H_{\texttt{core}}$ (see Claim 6.8), $H_{\texttt{core}}=H_{\texttt{init}}$ at the start of a phase, and $H_{\texttt{init}}$ is a $(B,(1-\epsilon)B)$-EDCS of $G_{\texttt{init}}$ (see Section 6.2). Thus, $H_{\texttt{init}}$ has maximum degree at most $B$ (see Definition 2.1), and hence $H_{\texttt{hilo}}$ also has maximum degree at most $B$. Accordingly, at the start of a phase, we can compute the subgraph $H_{\texttt{hilo}}$ in $O(nB)$ time. Beyond this $O(nB)$ term, the time taken to initialize $M_{\texttt{hilo}}$ and $M_{\texttt{base}}$ is dominated by: the time it takes to compute $M_{\texttt{most}}$, plus the time complexity of making one call to $\mathrm{Init}(.)$ and $|F^{\star}|\leq 6\delta n$ calls to $\mathrm{Augment}(.)$ (see Corollary 6.11 and part-(4) in Claim 6.12).
@@ 6.3.4 Maintaining the matchings $M_{\texttt{hilo}}$ and $M_{\texttt{base}}$ during the updates in a phase
    Excluding the time spent on initializing $M_{\texttt{base}}$ and $M_{\texttt{hilo}}$ (see Section 6.3.3), the total time spent on maintaining $M_{\texttt{base}}$ and $M_{\texttt{hilo}}$ within the phase is proportional to the time it takes to implement $\delta n$ calls to $\mathrm{Delete}(.)$ and $\mathrm{Augment}(.)$ in $H_{\texttt{hilo}}$ (see Lemma 5.1 and Claim 6.8).
@@
{mdframed}
[hidealllines=true, backgroundcolor=gray!20, leftmargin=0cm,innerleftmargin=0.4cm,innerrightmargin=0.4cm,innertopmargin=0.4cm,innerbottommargin=0.4cm,roundcorner=10pt] Both the matchings $M_{\texttt{base}}$ and $M_{\texttt{hilo}}$ remain unchanged if an edge gets inserted into the input graph $G=(V,E)$. Henceforth, consider an update that deletes an edge $e$ from $G$.
  • We first set $M_{\texttt{base}}\leftarrow M_{\texttt{base}}\setminus\{e\}$.
  • By part-(2c) of Claim 6.8, this update leads to at most one edge deletion (and zero edge insertions) in $H_{\texttt{hilo}}$. If no edge gets deleted from $H_{\texttt{hilo}}$, then the matching $M_{\texttt{hilo}}$ remains unchanged, and we proceed towards handling the next update in $G$.
@@
  • Otherwise, the same edge $e=(v,u)$, with $v\in V_{\texttt{hi}}$ and $u\in V_{\texttt{lo}}\cup V_{\texttt{a-lo}}$ (say), also gets deleted from $H_{\texttt{hilo}}$ (see Claim 6.8). In this case, we call $\mathrm{Delete}(v,u)$ in LPM (see Lemma 5.1). This call to $\mathrm{Delete}(.)$ leads to at most node in $V_{\texttt{sf}}$ getting unmatched under $M_{\texttt{hilo}}$, and if such a node exists, then it must be $v$. Under such an event (i.e., if $v\in V_{\texttt{sf}}\setminus V(M_{\texttt{hilo}})$), we call the subroutine $\mathrm{MatchViaAugment}(v)$ as described in Section 6.3.3. When this call terminates, we are ready to handle the next update in $G$.
Proof of Claim 6.13
@@
For part-(1), note that 6.10 holds before we start processing the very first update in the phase (see part-(1) of Claim 6.9). Now, consider any update in $G$ within the phase, and by inductive hypothesis, suppose that 6.10 holds just before this update. We have two cases.
@@
Case (i). The update in $G$ does not lead to any change in $H_{\texttt{hilo}}$. In this case, from (8) and Claim 6.8, it is easy to verify that 6.10 continues to hold after we finish processing this update.
@@
Case (ii). The concerned update in $G$ is the deletion of an edge $e=(v,u)$ that was already part of $H_{\texttt{hilo}}$. In this case, the edge $e$ gets deleted from $H_{\texttt{hilo}}$. We set $M_{\texttt{base}}\leftarrow M_{\texttt{base}}\setminus\{e\}$, and then call $\mathrm{Delete}(v,u)$ in LPM. From Lemma 5.1 and Claim 6.8, it is easy to verify that 6.10 holds when we finish the call to $\mathrm{Delete}(v,u)$. At this point, if we have $v\in V_{\texttt{sf}}\setminus V(M_{\texttt{hilo}})$, then finally we call the subroutine $\mathrm{MatchViaAugment}(v)$. Part-(1) of Claim 6.12 implies that 6.10 continues to hold when we finish the call to $\mathrm{MatchViaAugment}(v)$.
@@
For part-(2), note that before we start processing the very first update in the phase, every node in $V_{\texttt{sf}}$ is matched in $M_{\texttt{base}}$, and at most $18\delta n$ nodes in $V_{\texttt{med}}$ are free in $M_{\texttt{base}}$ (see part-(2) of Claim 6.9).
@@
Now, while handling an update within the phase, the (potential) call to $\mathrm{MatchViaAugment}(v)$ ensures that every node in $V_{\texttt{sf}}$ continues to remain matched in $M_{\texttt{base}}$ (see part-(2) of Claim 6.12), and it leads to at most two nodes changing their matched status in $M_{\texttt{base}}$ (see part-(3) of Claim 6.12). Finally, the steps we perform before making the (potential) call to $\mathrm{MatchViaAugment}(v)$ can also lead to at most two nodes changing their matched status in $M_{\texttt{base}}$. Taken together, this implies that while handling an update within the phase, there are at most four nodes in $V_{\texttt{med}}\subseteq V$ that change their status from “matched” to “free” in $M_{\texttt{base}}$.
@@
Part-(3) follows from part-(3) of Claim 6.12.
For part-(4), recall that a phase lasts for $\delta n$ updates in $G$, and the time spent on handling each such update is dominated by the (at most one) call we make to $\mathrm{AugmentViaMatch}(.)$. The bound on the total time complexity during a phase now follows from part-(4) of Claim 6.12.
6.3.5 Proof of Lemma 6.6
From part-(1), part-(2) and part-(3) of Claim 6.13, we infer that the matching $M_{\texttt{base}}$ satisfies Condition 6.4 and Condition 6.5. It now remains to analyze the amortized update time of the algorithm.
Suppose that we use the deterministic algorithms from Lemma 5.1 and Claim 6.7. First, observe that $E(H_{\texttt{hilo}})\subseteq E(H_{\texttt{core}})\subseteq E(H_{\texttt{init}})$, and $H_{\texttt{init}}$ is a $(B,(1-\epsilon)B)$-EDCS of $G_{\texttt{init}}$ (see Section 6.1 and Section 6.2). So, the maximum degree in $H_{\texttt{hilo}}$, denoted by $\Delta\left(H_{\texttt{hilo}}\right)$, is at most the maximum degree in $H_{\texttt{init}}$. This latter quantity, in turn, is at most $B$ (see Definition 2.1), and hence:
@@
Recall the values of $\gamma,X$ from (9) and that of $\delta,\epsilon$ from (6). Now, by part-(3) of Claim 6.9 and part-(4) of Claim 6.13, the time spent on maintaining the matching $M_{\texttt{base}}$ during a phase is at most:
@@
Next, suppose that we use the randomized algorithms from Lemma 5.1 and Claim 6.7. As before, recall the values of $\gamma,X$ from (9) and that of $\delta,\epsilon$ from (6). Now, by part-(3) of Claim 6.9 and part-(4) of Claim 6.13, the time spent on maintaining the matching $M_{\texttt{base}}$ during a phase is at most:

### v2

@@ 6.3.3 Initializing the matchings $M_{\texttt{hilo}}$ and $M_{\texttt{base}}$ at the start of a phase
Throughout this section, we focus on the snapshort of time at the start of a concerned phase, just before the first update (in that phase) arrives. Thus, for example, throughout this section we have $V_{\texttt{sf}}=V_{\texttt{hi}}$ and $E(H_{\texttt{init}})=E(H_{\texttt{core}})$ (see Definition˜6.2 and Corollary˜6.3). Furthermore, recall that $M(S)$ is the set of edges in a matching $M$ that are incident on $S$ (see Section˜2).
@@
    The total time spent to compute $M_{\texttt{hilo}}$ and $M_{\texttt{base}}$ is proportional to: $O(nB)$, plus the time it takes to compute $M_{\texttt{most}}$ (see Claim˜6.7), plus the time it takes to make one call to $\mathrm{Init}(.)$ and $6\delta n$ calls to $\mathrm{Augment}(.)$ in $H_{\texttt{hilo}}$ (see Lemma˜5.1 and part-(1) of Claim˜6.8).
Our algorithm will satisfy the following invariant, which implies part-(1) of Claim˜6.9.
@@
We devote the rest of this section to the proof of Claim˜6.9. Since $M_{\texttt{most}}\left(V_{\texttt{sf}}\right)\subseteq M_{\texttt{most}}\subseteq E(H_{\texttt{init}})=E(H_{\texttt{core}})$ (see Claim˜6.7), from (9) it follows that:
@@
Fix the parameters $\gamma,X$ as in (10), and recall Claim˜6.8. We will run the LPM data structure on $H_{\texttt{hilo}}$ (see Lemma˜5.1), and we will let $M_{\texttt{hilo}}$ be the matching $M$ maintained by LPM.
We first set $M_{\texttt{base}}\leftarrow M_{\texttt{most}}$, and then call $\mathrm{Init}(M_{\texttt{base}}(V_{\texttt{sf}}))$ in LPM. [^8] This is a valid operation, since $M_{\texttt{most}}(V_{\texttt{sf}})\subseteq E(H_{\texttt{hilo}})$ according to (12).
@@
At the end of the above call to $\mathrm{Init}(.)$, at most $6\delta n$ nodes from $V_{\texttt{sf}}\cup V_{\texttt{med}}$ are free under $M_{\texttt{base}}$, and ˜6.10 is satisfied.
@@
Follows from Claims˜6.7 and 6.8 and Lemma˜5.1. ∎
Let $F^{\star}=V_{\texttt{sf}}\setminus V(M_{\texttt{hilo}})$ denote the set of safe high nodes that are free under $M_{\texttt{hilo}}$ at this point. By Corollary˜6.11 and Claim˜6.8, we have $|F^{\star}|\leq 6\delta n$, and $\mathrm{deg}_{H_{\texttt{hilo}}}(v)\geq X$ for all $v\in F^{\star}$. Next, for each node $v\in F^{\star}$, we call the subroutine $\mathrm{MatchViaAugment}(v)$, as described below. This concludes our initialization of the matching $M_{\texttt{base}}$ at the start of a phase.
@@
We call this subroutine only if $v$ is currently unmatched in $M_{\texttt{hilo}}$ and $v\in V_{\texttt{sf}}$, so that $v$ lies on the “left” with $\mathrm{deg}_{\texttt{hilo}}(v)\geq X$ (see Claim˜6.8). It performs the following operations. Make a call to $\mathrm{Augment}(v)$ in LPM (see Lemma˜5.1). This step matches $v$ under $M_{\texttt{hilo}}$ by applying an augmenting path, adds exactly one node (say, $u_{v}$) besides $v$ to the set $V(M_{\texttt{hilo}})$, and does not remove any node from the set $V(M_{\texttt{hilo}})$. Since $v$ lies to the “left” of the bipartite graph $H_{\texttt{hilo}}$, it follows that $u_{v}$ lies to the “right” of $H_{\texttt{hilo}}$ and hence $u_{v}\in V_{\texttt{lo}}\cup V_{\texttt{a-lo}}$ (see Claim˜6.8). Let $M^{\texttt{old}}_{v}$ (resp. $M^{\texttt{new}}_{v}$) denote the set of edges that get removed from (resp. added to) the matching $M_{\texttt{hilo}}$ as we apply this augmenting path. Set $M_{\texttt{base}}\leftarrow M_{\texttt{base}}\cup M^{\texttt{new}}_{v}\setminus M^{\texttt{old}}_{v}$. If $u_{v}$ was previously matched under $M_{\texttt{base}}$ (say, via an edge $(u_{v},u^{\prime}_{v})\in M_{\texttt{base}}$), then set $M_{\texttt{base}}\leftarrow M_{\texttt{base}}\setminus\{(u_{v},u^{\prime}_{v})\}$.
@@
For part-(2), the node $v$ gets added to the set $V(M_{\texttt{base}})$ because of the call to $\mathrm{Augment}(v)$ in LPM (see Lemma˜5.1), and because we ensure that $M_{\texttt{base}}\supseteq M_{\texttt{hilo}}$ (see part-(1) above). Furthermore, $u^{\prime}_{v}$ is the only node that can potentially get removed from $V(M_{\texttt{base}})$. Just before the call to $\mathrm{MatchViaAugment}(v)$, we have $(u_{v},u^{\prime}_{v})\notin M_{\texttt{hilo}}$ (otherwise the augmenting path we apply will not end at $u_{v}$) and $M_{\texttt{hilo}}=M_{\texttt{base}}\left(V_{\texttt{sf}}\right)$ (due to ˜6.10). Thus, it follows that $u^{\prime}_{v}\notin V_{\texttt{sf}}$.
@@
Proof of Claim˜6.9
Part-(1) follows from ˜6.10, Corollary˜6.11 and part-(1) of Claim˜6.12.
For part-(2), note that $\left|\left(V_{\texttt{sf}}\cup V_{\texttt{med}}\right)\setminus V(M_{\texttt{base}})\right|\leq 6\delta n$ at the end of the call to $\mathrm{Init}(.)$ in LPM (see Corollary˜6.11). Subsequently, for each node $v\in F^{\star}=V_{\texttt{sf}}\setminus V(M_{\texttt{hilo}})$, we call $\mathrm{MatchViaAugment}(v)$, which leads to:
  • the node $v$ getting added to $V(M_{\texttt{base}})$ and no node already in $V_{\texttt{sf}}\cap V(M_{\texttt{base}})$ getting deleted from $V(M_{\texttt{base}})$ (see part-(2) of Claim˜6.12), and
  • at most two nodes in $V_{\texttt{med}}$ getting deleted from $V(M_{\texttt{base}})$ (see part-(3) of Claim˜6.12).
@@
For part-(3), note that $H_{\texttt{hilo}}$ is a subgraph of $H_{\texttt{core}}$ (see Claim˜6.8), $H_{\texttt{core}}=H_{\texttt{init}}$ at the start of a phase, and $H_{\texttt{init}}$ is a $(B,(1-\epsilon)B)$-EDCS of $G_{\texttt{init}}$ (see Section˜6.2). Thus, $H_{\texttt{init}}$ has maximum degree at most $B$ (see Definition˜2.1), and hence $H_{\texttt{hilo}}$ also has maximum degree at most $B$. Accordingly, at the start of a phase, we can compute the subgraph $H_{\texttt{hilo}}$ in $O(nB)$ time. Beyond this $O(nB)$ term, the time taken to initialize $M_{\texttt{hilo}}$ and $M_{\texttt{base}}$ is dominated by: the time it takes to compute $M_{\texttt{most}}$, plus the time complexity of making one call to $\mathrm{Init}(.)$ and $|F^{\star}|\leq 6\delta n$ calls to $\mathrm{Augment}(.)$ (see Corollary˜6.11 and part-(4) in Claim˜6.12).
@@ 6.3.4 Maintaining the matchings $M_{\texttt{hilo}}$ and $M_{\texttt{base}}$ during the updates in a phase
    Excluding the time spent on initializing $M_{\texttt{base}}$ and $M_{\texttt{hilo}}$ (see Section˜6.3.3), the total time spent on maintaining $M_{\texttt{base}}$ and $M_{\texttt{hilo}}$ within the phase is proportional to the time it takes to implement $\delta n$ calls to $\mathrm{Delete}(.)$ and $\mathrm{Augment}(.)$ in $H_{\texttt{hilo}}$ (see Lemma˜5.1 and Claim˜6.8).
@@
Both the matchings $M_{\texttt{base}}$ and $M_{\texttt{hilo}}$ remain unchanged if an edge gets inserted into the input graph $G=(V,E)$. Henceforth, consider an update that deletes an edge $e$ from $G$. We first set $M_{\texttt{base}}\leftarrow M_{\texttt{base}}\setminus\{e\}$. By part-(2c) of Claim˜6.8, this update leads to at most one edge deletion (and zero edge insertions) in $H_{\texttt{hilo}}$. If no edge gets deleted from $H_{\texttt{hilo}}$, then the matching $M_{\texttt{hilo}}$ remains unchanged, and we proceed towards handling the next update in $G$. Otherwise, the same edge $e=(v,u)$, with $v\in V_{\texttt{hi}}$ and $u\in V_{\texttt{lo}}\cup V_{\texttt{a-lo}}$ (say), also gets deleted from $H_{\texttt{hilo}}$ (see Claim˜6.8). In this case, we call $\mathrm{Delete}(v,u)$ in LPM (see Lemma˜5.1). This call to $\mathrm{Delete}(.)$ leads to at most node in $V_{\texttt{sf}}$ getting unmatched under $M_{\texttt{hilo}}$, and if such a node exists, then it must be $v$. Under such an event (i.e., if $v\in V_{\texttt{sf}}\setminus V(M_{\texttt{hilo}})$), we call the subroutine $\mathrm{MatchViaAugment}(v)$ as described in Section˜6.3.3. When this call terminates, we are ready to handle the next update in $G$.
Proof of Claim˜6.13
@@
For part-(1), note that ˜6.10 holds before we start processing the very first update in the phase (see part-(1) of Claim˜6.9). Now, consider any update in $G$ within the phase, and by inductive hypothesis, suppose that ˜6.10 holds just before this update. We have two cases.
@@
Case (i). The update in $G$ does not lead to any change in $H_{\texttt{hilo}}$. In this case, from (9) and Claim˜6.8, it is easy to verify that ˜6.10 continues to hold after we finish processing this update.
@@
Case (ii). The concerned update in $G$ is the deletion of an edge $e=(v,u)$ that was already part of $H_{\texttt{hilo}}$. In this case, the edge $e$ gets deleted from $H_{\texttt{hilo}}$. We set $M_{\texttt{base}}\leftarrow M_{\texttt{base}}\setminus\{e\}$, and then call $\mathrm{Delete}(v,u)$ in LPM. From Lemma˜5.1 and Claim˜6.8, it is easy to verify that ˜6.10 holds when we finish the call to $\mathrm{Delete}(v,u)$. At this point, if we have $v\in V_{\texttt{sf}}\setminus V(M_{\texttt{hilo}})$, then finally we call the subroutine $\mathrm{MatchViaAugment}(v)$. Part-(1) of Claim˜6.12 implies that ˜6.10 continues to hold when we finish the call to $\mathrm{MatchViaAugment}(v)$.
@@
For part-(2), note that before we start processing the very first update in the phase, every node in $V_{\texttt{sf}}$ is matched in $M_{\texttt{base}}$, and at most $18\delta n$ nodes in $V_{\texttt{med}}$ are free in $M_{\texttt{base}}$ (see part-(2) of Claim˜6.9).
@@
Now, while handling an update within the phase, the (potential) call to $\mathrm{MatchViaAugment}(v)$ ensures that every node in $V_{\texttt{sf}}$ continues to remain matched in $M_{\texttt{base}}$ (see part-(2) of Claim˜6.12), and it leads to at most two nodes changing their matched status in $M_{\texttt{base}}$ (see part-(3) of Claim˜6.12). Finally, the steps we perform before making the (potential) call to $\mathrm{MatchViaAugment}(v)$ can also lead to at most two nodes changing their matched status in $M_{\texttt{base}}$. Taken together, this implies that while handling an update within the phase, there are at most four nodes in $V_{\texttt{med}}\subseteq V$ that change their status from “matched” to “free” in $M_{\texttt{base}}$.
@@
Part-(3) follows from part-(3) of Claim˜6.12.
For part-(4), recall that a phase lasts for $\delta n$ updates in $G$, and the time spent on handling each such update is dominated by the (at most one) call we make to $\mathrm{AugmentViaMatch}(.)$. The bound on the total time complexity during a phase now follows from part-(4) of Claim˜6.12.
6.3.5 Proof of Lemma˜6.6
From part-(1), part-(2) and part-(3) of Claim˜6.13, we infer that the matching $M_{\texttt{base}}$ satisfies Condition˜6.4 and Condition˜6.5. It now remains to analyze the amortized update time of the algorithm.
Suppose that we use the deterministic algorithms from Lemma˜5.1 and Claim˜6.7. First, observe that $E(H_{\texttt{hilo}})\subseteq E(H_{\texttt{core}})\subseteq E(H_{\texttt{init}})$, and $H_{\texttt{init}}$ is a $(B,(1-\epsilon)B)$-EDCS of $G_{\texttt{init}}$ (see Section˜6.1 and Section˜6.2). So, the maximum degree in $H_{\texttt{hilo}}$, denoted by $\Delta\left(H_{\texttt{hilo}}\right)$, is at most the maximum degree in $H_{\texttt{init}}$. This latter quantity, in turn, is at most $B$ (see Definition˜2.1), and hence:
@@
Recall the values of $\gamma,X$ from (10) and that of $\delta,\epsilon$ from (7). Now, by part-(3) of Claim˜6.9 and part-(4) of Claim˜6.13, the time spent on maintaining the matching $M_{\texttt{base}}$ during a phase is at most:
@@
Next, suppose that we use the randomized algorithms from Lemma˜5.1 and Claim˜6.7. As before, recall the values of $\gamma,X$ from (10) and that of $\delta,\epsilon$ from (7). Now, by part-(3) of Claim˜6.9 and part-(4) of Claim˜6.13, the time spent on maintaining the matching $M_{\texttt{base}}$ during a phase is at most: